# Transientangelo: Few-Viewpoint Surface Reconstruction Using Single-Photon Lidar

### v1

@@ 5.2 Captured Results
Figure 4 shows the recovered meshes for our method and the baselines on the food, carving, cinema scenes for two, three, and five training viewpoints respectively (see the supplement for qualitative results for all scenes). As in TransientNeRF [29], we mask out the background of the transients for the captured dataset. Similarly to the simulated dataset, the meshes recovered from the proposed method are both smoother and contain fewer artifacts than the baseline methods.
We report metrics averaged across six captured scenes in Table 2. Due to a lack of ground-truth meshes, we are not able to report Chamfer distances and the depth L1 metric is calculated by comparing to depth maps estimated from log-matched filtering of transients from held out views. Our quantitative results show an improvement over TransientNeRF across all image and depth metrics. Overall, our method remains robust across various photon levels, whereas TransientNeRF’s image rendering quality deteriorates rapidly as the mean photon count decreases.
Figure 5 shows the results of training our method and TransientNeRF [29] on transients with an average of $50$ and $300$ photons per occupied pixel. At these photon levels, TransientNeRF rendered transients’ intensities are weaker than the measured ones. On the other hand, due to the reflectivity loss (see supplement), our method recovers more accurate transients.
@@ 5.3 Ablation Study
Table 3 ablates the effect of the regularization terms on the simulated Lego scene. The results show that all the added regularizers are vital to the performance; however, the weight-variance regularizer has the greatest effect. Applying this regularizer to the rendered unseen viewpoints is especially important in the few-viewpoint setting.
@@ 6 Discussion
There are multiple limitations of our method that could be addressed in future work. For example, computational efficiency could potentially be improved by developing new time-resolved rendering frameworks based on Gaussian splatting [16]. While the main focus of our work is scene reconstruction using the direct component of light, transients carry much richer scene information, including higher order bounces of light. Future work could investigate such techniques for systems that infer further scene properties such as reflectance, albedo, or other material properties.

### v4

@@ 5.2 Captured Results
Figure 4 shows the recovered meshes for our method and the baselines on the food, carving, cinema scenes for two, three, and five training viewpoints respectively (see the supplement for qualitative results for all scenes). As in TransientNeRF [30], we mask out the background of the transients for the captured dataset. Similarly to the simulated dataset, the meshes recovered from the proposed method are more accurate than the baseline methods.
We report metrics averaged across six captured scenes in Table 2. Due to a lack of ground-truth meshes, we are not able to report Chamfer distances, and the depth L1 metric is calculated by comparing depth maps estimated from log-matched filtering of transients from held-out views. Our quantitative results show an improvement over TransientNeRF across all image and depth metrics. Overall, our method remains robust across various photon levels, whereas TransientNeRF’s image rendering quality deteriorates rapidly as the mean photon count decreases.
Figure 5 shows the results of training our method and TransientNeRF [30] on transients with an average of $50$ and $300$ photons per occupied pixel. At these photon levels, using TransientNeRF results in rendered transients that underestimate the measured intensities compared to our method.
@@ 5.3 Ablation Study
Table 3 evaluates the effect of ablating the regularization terms on the simulated Lego scene. The results indicate that all the added regularizers are crucial for performance; however, $\mathcal{L}_{\text{ref}}$ and $\mathcal{L}_{\text{weight\_var}}$ are the most essential. When ablating the reflectivity loss (row 4), we deviated from the default settings of $\mathcal{L}_{\text{eik}}$ and $\mathcal{L}_{\text{weight\_var}}$, as indicated in the table, to improve convergence (see details in the supplement).
@@ 6 Discussion
In future work, we hope to address certain limitations of our method. For example, computational efficiency could potentially be improved by developing new time-resolved rendering frameworks based on Gaussian splatting [16]. Additionally, while the main focus of our work is scene reconstruction using the direct component of light, transients carry much richer scene information, including higher order bounces of light, future work could exploit indirect light transport to infer scene properties such as geometry, reflectance, and material characteristics.
7 Acknowledgments
DBL acknowledges support of the Canada Foundation for Innovation, the Ontario Research Fund, and NSERC under the RGPIN, RTI, and Alliance programs.